OpenStudy (anonymous):

What is the slope of the line 5x - 2y = 10?

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

This is because if you simplify the equation it will look like this y=5/2x-5 which is in the form y=mx+b with m equaling the slope

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

so 5/2 is the answer right?